Status Report

NASA Space Station On-Orbit Status 23 May 2003

By SpaceRef Editor
May 23, 2003
Filed under , ,
NASA Space Station On-Orbit Status 23 May 2003

All ISS systems continue to function nominally, except as noted previously or below.

CDR Yuri Malenchenko completed the regular weekly checkout of the IP-1 air flow sensors which are installed in the various RS (Russian segment) hatchways and the hatchway between FGB and Node.

Malenchenko also performed an inventory (IMS) audit of personal hygiene items (SGO), including garments, sleeping bag liners, etc., as well as microgravity countermeasures items, such as Penguin suits, remaining from previous crews.  He prepared them for stowage for subsequent crews or disposal per instructions uplinked overnight.

Later, Yuri Malenchenko undertook another weekly inventory audit of the available CWCs (collapsible water containers) and their condition.  [Last time done: 5/9.  [Expedition 7 crew is observing drinking water rationing (without detriment to health maintenance, for example, no renal stone risk), but besides potable water enough “technical” water needs to be kept in reserve in case of unexpected recharge requirements for the ITCS (internal thermal control system).  In order to maintain good chemistry in the cooling loops, technical water for ITCS does not contain the minerals that are added to potable water for its safe use.]

FE/SO Edward Lu supported another session with the InSPACE (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions) experiment in the MSG (Microgravity Science Glovebox).  This was run #11 for the present coil assembly (CA005).  [The test utilized the lowest current field strength, with amplitude set at 0.875 amps for the entire run.  Initial frequency was 5.0 Hz, later switched to 2.0 Hz for Lu’s second monitor period.  The procedures were updated to allow for removal of the not-needed MSG glove for adjusting the potentiometer.  The MSG was later powered down again.]

Yuri conducted a checkup of the stowage conditions for defective Vozdukh BVK vacuum valve packages and BVN vacuum pump units, with an audit of BVK units stowed and currently in use.  [RSC-Energia wants to ensure that the BVN pump inlets and the BVK valve fittings are properly capped and their connectors covered for protection during stowage.  These units are potentially recyclable after return to Earth.]

Ed performed the periodic transfer of accumulated data files from the CMS HRM (crew medical systems/heart rate monitor) to the MEC (medical equipment computer) for downlink, then deleted them on the HRM. 

Later in the day, he transferred data files from the physical exercise equipment (TVIS and RED) to the MEC via memory card, also for downlink via OCA comm.

Malenchenko conducted the daily routine maintenance of the SM’s SOZh life support system (including ASU toilet facilities) and prepared the daily IMS inventory delta file, while Ed Lu did the regular status checkup of the autonomous Lab payload (PCG-STES010, SAMS, MAMS).

For the third time, Lu completed his weekly task of filling out the FFQ (food frequency questionnaire), which keeps a log of his nutritional intake over time on special MEC software.

The crew conducted the regular physical exercise protocol on the treadmill, followed by workout sessions on the RED (resistive exercise device) and the Russian VELO bike with load trainer.

Ed and Yuri also had their weekly conferences with the JSC Astronaut Office (Kent Rominger) and ISS Program management.  Later in the day, their schedule called for the weekly teleconference with the Flight Director at MCC-H.

The newly cropped-up discrepancy between CO2 partial pressure (ppCO2) readings in the Lab module and the SM (see yesterday’s status) has been resolved.  Using the U.S. CDMK (carbon dioxide monitor kit), a value of 4.1 mmHg CO2 was measured last night in the SM and 4.2 mmHg in the Lab.  This indicates that ventilation and mixing across the station is good and that the RS CO2 measurements were erroneous..

As every day since last Tuesday (5/20), TsUP performed another test correction of the SM’s BINS strap-down navigation/guidance system via S-band comm, using Zvezda’s solar sensor (SD251K) and infrared horizon sensor (IKV256K).  [The test correction in the current LVLH -YVV attitude does not require free drift or attitude change.  BINS is a platform-less (or, in NASA parlance, “strap-down”) inertial navigation system (INS), which uses a complex mathematical model of spatial relationships (axes, rotations, angular velocities and accelerations) instead of the gyro-stabilized and gimbaled platform of a more conventional INS.  To check and correct the unavoidable drift of the internal reference orientation, the standard procedure was to take optical sightings of reference stars; attitude rates are obtained from the GIVUS high-accuracy state sensors.  Today, periodic updates of the BINS come from the U.S. GPS (global positioning system).]

Working from the Russian task list, Malenchenko was to use some of his free time for the Russian Starmail Earth photography experiment.  [The Nikon D1 digital camera with AF-S Nikkor lens is used, and some frames downlinked earlier showed smudges looking like “soft focus spots”.  Yuri was requested to investigate possible causes and to select the cleanest window for Earth photography.  The ground is looking into methods for cleaning the camera’s highly sensitive CCD (charged-couple device) matrix, perhaps by softly blowing particulates off the mirror with air.]

A new item added to the NASA “job jar” task list is an evaluation of closeout panels in the Airlock (AL) “Quest”, to document broken fasteners, current panel configuration and any configuration issues.  [On previous flights, crews have noted difficulty with overlapping panels, broken fasteners, etc. in the AL, but no specifics regarding the extent of the problem have been documented as yet.  The assessment, which is needed on the ground to assess kick-load and panel grounding concerns, must be completed by Monday (5/26) to allow time for assessing the results before the EMU don/doff test on Wednesday.]

Also added to the task list was a checkup on two U.S./RS hoses on board.  [The US/RS hoses, for transferring water from U.S. CWCs to Russian EDV soft water containers, or from CWCs to the Elektron’s CFU (condensate feed unit), have been stored in the SM. Once wetted, these hoses have a lifetime of 5 years, and the ground needs verification on whether their use has been started yet.] 

The crew was thanked for their great support of yesterday’s PAO/educational event with the Adler Planetarium in Chicago, IL.  [The Planetarium’s Educational Director is Ms. Michelle Nichols, not Nichelle N., as mistakenly reported yesterday.]

Today’s CEO (crew earth observations) targets, now no longer restricted by the Lab science window ruled off-limit due to flight attitude, were Lake Eyre, Australia (recent Increment 6 images showed widespread filling of small lakes around Eyre, indicating unusual local rains about 3-4 months ago.  The west lobe of Lake Eyre itself is now slowly filling as water drains through sand dunes to the low point of the basin.  An update on this important sensor of central Australian climate is requested), Sao Paulo, Brazil (nadir pass), Rio de Janeiro, Brazil (a touch right of track), Saharan dust, West Africa (Dynamic event. Major dust storm winding down: looking left for the dust-cloud mass over the Atlantic Ocean and trying to include Cape Verde Islands in the view to allow for orientation), Chao Lava, Chile (visible again today on an ascending track, this major lava flow appears as a black wrinkled mass in the desert of the high Andes [this site was seen on yesterday’s descending track]), Central Amazon River Basin (Dynamic event.  Sunglint should have illuminated a large area on the unusually clear south side of the Amazon River.  River patterns in the region are very poorly mapped with respect to the real complexity of the drainage.  Looking left on the nearer side of the glint disc), Lower Amazon River Basin (looking right for detailed views of the shorelines and islands of the estuary of the Amazon River mouth), Dakar, Senegal(night target; nadir pass), St. Louis, Missouri (night target; nadir pass), Lahore, Pakistan (night target; nadir pass), Osaka, Japan (night target; nadir pass), and Rome, Italy (night target; nadir pass).
CEO images can be viewed at the websites and

U.S. and Russian Segment Status (as of 1:45pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (16 Amp mode).  Vozdukh CO2 scrubber is On (manual mode 3).  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is operating.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 27.4; ppO2 (mmHg) — data invalid; ppCO2 (mmHg) — data invalid.
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 22.3.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — 21.3.
  • Node:  Pressure (mmHg) — 744.52; temperature (deg C) — 23.4 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 746.33; temperature (deg C) — 22.5; ppO2 (mmHg) — 172.7; ppCO2 (mmHg) — 4.8.
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 746.43; temperature (deg C) — 28.2; shell heater temp (deg C) — 24.7, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 21.8.
  • PMA-2:  Shell heater temp (deg C) — 13.6.

(n/a = data not available)

Propulsion System (PS):

  • Total propellant load available: 3573 kg (7877 lb) as of 5/22 [SM(774) + FGB(2547) + Progress(252)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both “directed” (non-solar tracking) mode (“sun slicer” position, for drag reduction).
  • SM batteries:  Batteries #4 and #8 are disconnected (failed); all other batteries (6) are in “Partial Charge” mode (battery #2 is exhibiting degraded performance).
  • FGB batteries:  Battery #4 is disconnected; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Discharge mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-1 is operating; INT-2 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is Off (standby); PL-1 MDM is On
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational. Being upgraded.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational. Being upgraded.

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-2 (GPS)
  • Angular rate source — RGA-2

Flight Attitude:

  • VLH -YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, -y-axis in velocity vector [yaw: -90 deg, pitch: -6.8 deg, roll: -3.4 deg]), with CMG/TA (thruster assist) Momentum Management.
  • Solar Beta angle:  71.5 deg (magnitude decreasing).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally [IAC-1 (internal audio controller #1) being analyzed after self-test error.  IAC-2 is prime.]
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1 with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched at WS4, with KA power. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit  (as of this morning, 8:11am EDT [= epoch]):

  • Mean altitude — 389.2 km
  • Apogee — 393.3 km
  • Perigee — 385.2 km
  • Period — 92.34 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0005975
  • Orbits per 24-hr. day — 15.59
  • Solar Beta Angle — 71.5 deg (magnitude decreasing)
  • Mean altitude loss in last 24 hours — 75 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 25715
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.